# Morita equivalence of Nijenhuis structures

### v1

@@ 4.4 Remarks on hierarchies of Morita equivalent Dirac-Nijenhuis structures
is a Morita equivalence of the Dirac-Nijenhuis structures $\big(M_{i},L_{i}^{(0,n)},r_{i}\big)$, where $\varpi_{n}^{\flat}\coloneqq\varpi^{\flat}\circ J^{n}$.
We prove first for $n=1$. The actions of $L_{1}$ and $L_{2}$ on $P$ are given by $(\mathrm{d}\mu_{i}(v_{i}),\alpha_{i})\in\Gamma(L_{i})\mapsto v_{i}\in\mathfrak{X}(P)$ where $\varpi^{\flat}(v_{i})=(\mathrm{d}\mu_{i})^{*}\alpha_{i}$. We need to prove that
[…]
Indeed, this is the case since
$\varpi_{1}^{\flat}(v_{i})$ $=\varpi^{\flat}(J(v_{i}))$
$=J^{*}\varpi^{\flat}(v_{i})$
$=J^{*}(\mathrm{d}\mu_{i})^{*}\alpha_{i}$
$=(\mathrm{d}\mu_{i}\circ J)^{*}\alpha_{i}$
$=(r_{i}\circ\mathrm{d}\mu_{i})^{*}\alpha_{i}$
$=(\mathrm{d}\mu_{i})^{*}r_{i}^{*}\alpha_{i}.$
This computation shows that $\mu_{1}$ and $\mu_{2}$ are forward Dirac maps. The induced actions
$\mathscr{a}_{1}^{(0,1)}:\Gamma\big(L_{1}^{(0,1)}\big)\to\mathfrak{X}(P)\qquad\text{and}\qquad\mathscr{a}_{2}^{(0,1)}:\Gamma\big(L_{2}^{(0,1)}\big)\to\mathfrak{X}(P)$
send sections of the form $\xi=(\mathrm{id}_{TM_{1}},r_{1}^{*})(\xi_{0})\in\Gamma\big(L_{1}^{(0,1)}\big)$ and $\zeta=(\mathrm{id}_{TM_{2}},r_{2}^{*})(\zeta_{0})\in\Gamma\big(L_{2}^{(0,1)}\big)$, where $\xi_{0}=(\mathrm{d}\mu_{1}(v_{1}),\alpha_{1})$ and $\zeta_{0}=(\mathrm{d}\mu_{2}(v_{2}),\alpha_{2})$ to the vector fields $v_{1}$ and $v_{2}$. In other words,
$\mathscr{a}_{1}^{(0,1)}(\xi)=\mathscr{a}_{1}(\xi_{0})\qquad\text{and}\qquad\mathscr{a}_{2}^{(0,1)}(\zeta)=\mathscr{a}_{2}(\zeta_{0}).$ (55)
@@
Therefore it follows immediately that these actions satisfy all conditions of the Morita equivalence of Dirac structures (Definition 4.11) and of infinitesimal Nijenhuis structures (Definition 3.12). Moreover, as we noted before, the $2$-form $\varpi_{1}$ is compatible with $J$ and therefore we have constructed a Dirac-Nijenhuis Morita equivalence.
Let us suppose that $\big(L_{1}^{(0,n-1)},r_{1}\big)$ and $\big(L_{2}^{(0,n-1)},r_{2}\big)$ are Morita equivalent via $(P,\varpi_{n-1},J)$ and proceed by induction. The actions of $L_{1}^{(0,n-1)}$ and $L_{2}^{(0,n-1)}$ are given by $\mathscr{a}_{1}^{(0,n-1)}\left((\mathrm{id}_{TM_{1}},(r_{1}^{*})^{n-1})(\xi_{0})\right)=\mathscr{a}_{1}(\xi_{0})$ and $\mathscr{a}_{2}^{(0,n-1)}\left((\mathrm{id}_{TM_{2}},(r_{2}^{*})^{n-1})(\zeta_{0})\right)=\mathscr{a}_{2}(\zeta_{0})$. The same computation as above shows that $\varpi_{n}^{\flat}(v_{i})=(\mathrm{d}\mu_{i})^{*}(r^{n}_{i})^{*}\alpha_{i}$ for $i=1,2$ and
$\mathscr{a}_{1}^{(0,n)}(\xi)=\mathscr{a}_{1}(\xi_{0})\qquad\text{and}\qquad\mathscr{a}_{2}^{(0,n)}(\zeta)=\mathscr{a}_{2}(\zeta_{0}),$ (56)
which finishes the proof since $\varpi_{n}$ and $J$ are compatible. ∎
@@
Let $(M_{1},\pi_{1},r_{1})$ and $(M_{2},\pi_{2},r_{2})$ be Morita equivalent Poisson-Nijenhuis structures via a bibundle
@@
If $J\in\Omega^{1}(P,TP)$ is an invertible Nijenhuis $(1,1)$-tensor, then the Poisson-Nijenhuis structures $(M_{1},\pi_{1}^{n},r_{1})$ and $(M_{2},\pi_{2}^{n},r_{2})$ are (weakly) Morita equivalent for all $n\in\mathbb{N}$.
@@
To obtain a (strong) Morita equivalence, this proposition requires the assumption that the infinitesimal actions associated with the hierarchy are complete. These actions are defined by deforming the original actions $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ with powers of $J$ (as detailed in the construction below) and are not automatically complete. Without this completeness assumption, the structures are only weakly Morita equivalent in the sense of Ginzburg [20].
@@
Since $(\varpi,J)$ is a compatible pair and $\varpi$ is symplectic, $(P,\pi_{\varpi},J)$ is a Poisson-Nijenhuis manifold, where $\pi_{\varpi}=-(\varpi^{\flat})^{-1}$. Consequently, there exists a hierarchy of Poisson-Nijenhuis structures $(\pi_{\varpi}^{n},J)$ on $P$. Note that each $\pi_{\varpi}^{n}$ is a non-degenerate Poisson structure. Indeed, $J^{n}$ is invertible since $J$ is invertible; furthermore, $\pi_{\varpi}^{\sharp}$ is an isomorphism by the non-degeneracy of $\varpi$. It follows that the composition $\pi_{\varpi}^{n\sharp}=J^{n}\circ\pi_{\varpi}^{\sharp}$ is invertible. We define the symplectic forms $\varpi^{n}\in\Omega^{2}(P)$ by $\varpi^{n\flat}\coloneqq-(\pi_{\varpi}^{n\sharp})^{-1}$, i.e. $\varpi^{n\flat}=\varpi^{\flat}\circ J^{-n}$. Finally, the bibundle for the $n$-th level of the hierarchy is given by the bibundle
$(M_{1},\pi_{1}^{n},r_{1})\xleftarrow{\quad\mu_{1}\quad}(P,\varpi^{n},J)\xrightarrow{\quad\mu_{2}\quad}(M_{2},\pi_{2}^{n},r_{2}).$
The infinitesimal actions are defined by deforming the original actions with the tensor field $J^{n}$:
$\mathscr{a}^{(n,0)}_{1}\left(\pi^{n\ \sharp}_{1}(\alpha),\alpha\right)=J^{n}(u),\qquad\text{and}\qquad\mathscr{a}^{(n,0)}_{2}\left(\pi^{n\ \sharp}_{2}(\beta),\beta\right)=J^{n}(v),$
where $u,v\in\mathfrak{X}(P)$ are such that $\mathrm{d}\mu_{1}(u)=\pi^{\sharp}_{1}(\alpha)$ and $\mathrm{d}\mu_{2}(v)=\pi^{\sharp}_{2}(\beta)$. Provided these actions are complete, this bibundle establishes the Morita equivalence at the $n$-th level.
Modular vector field of a Poisson-Nijenhuis manifold.
@@
We now comment on the cohomological consequences arising from the hierarchy of Morita equivalent Poisson–Nijenhuis structures. Henceforth, we consider an infinitesimal Morita bibundle of Poisson–Nijenhuis structures
$(M_{1},\pi_{1},r_{1})\xleftarrow{\quad\mu_{1}\quad}(P,\varpi,J)\xrightarrow{\quad\mu_{2}\quad}(M_{2},\pi_{2},r_{2}),$
where $J\in\Omega^{1}(P,TP)$ is an invertible Nijenhuis tensor field, which we call non-degenerate infinitesimal Morita bibundle. Thus, the hypotheses of Proposition 4.21 are satisfied, and there exists a hierarchy of infinitesimal bibundles
$(M_{1},\pi_{1}^{n},r_{1})\xleftarrow{\ \mu_{1}\ }(P,\varpi^{n},J)\xrightarrow{\ \mu_{2}\ }(M_{2},\pi_{2}^{n},r_{2}).$
Although the infinitesimal actions in this hierarchy are not necessarily all complete, and therefore do not define Morita bundles in our sense, they still constitute weak Morita equivalences [20] and the present results remain valid.
@@
Every Poisson manifold $(M_{i},\pi_{i})$ induces a differential $\mathrm{d}_{\pi_{i}}:\mathfrak{X}^{k}(M_{i})\to\mathfrak{X}^{k+1}(M_{i})$. The cohomology $H^{k}_{\pi}(M_{i})$ of the complex $(\mathfrak{X}^{\bullet}(M_{i}),\mathrm{d}_{\pi_{i}})$ is called the Poisson cohomology of $(M_{i},\pi_{i})$. It was shown in [19] that Morita equivalent Poisson manifolds have isomorphic first Poisson cohomology groups; that is, $H^{1}_{\pi_{1}}(M_{1})\cong H^{1}_{\pi_{2}}(M_{2})$. In our setting, we thus obtain a hierarchy of isomorphic first Poisson cohomologies. For every $n\in\mathbb{N}$, it follows that
$H^{1}_{\pi_{1}^{n}}(M_{1})\cong H^{1}_{\pi_{2}^{n}}(M_{2}).$
@@
We recall the modular vector field of a Poisson manifold. Assume for the moment that $M_{i}$ is orientable. The divergence of a vector field $v\in\mathfrak{X}(M_{i})$ with respect to a volume form $\nu_{i}$ is the unique function $\mathrm{div}_{\nu_{i}}(v)\in C^{\infty}(M_{i})$ such that $\mathscr{L}_{v}(\nu_{i})=\mathrm{div}_{\nu_{i}}(v)\nu_{i}$. For an arbitrary function $f\in C^{\infty}(M_{i})$, let $X_{f}=\pi_{i}^{\sharp}(\mathrm{d}f)$ denote its Hamiltonian vector field. The modular vector field of $(M_{i},\pi_{i})$ is the Poisson vector field $X_{\pi_{i}}^{\nu_{i}}\in\mathfrak{X}(M_{i})$ defined by
$X_{\pi_{i}}^{\nu_{i}}(f)\coloneqq\mathrm{div}_{\nu_{i}}\left(X_{f}\right).$
@@
This vector field depends on the choice of the volume, but the cohomology class $\left[X_{\pi_{i}}^{\nu_{i}}\right]\in H^{1}_{\pi_{1}^{n}}(M)$ does not. This $\left[X_{\pi_{i}}^{\nu_{i}}\right]$ is called the modular class of the Poisson manifold $(M_{i},\pi_{i})$. By replacing volume forms with smooth densities in the definition of the modular vector field, we may drop the orientability requirement on the manifold $M$. Because these classes correspond under the isomorphism of the first Poisson cohomologies [12, Corollary 7], we obtain that any Morita equivalence via a non-degenerate infinitesimal Morita bibundle $(P,\varpi,J)$ between Poisson-Nijenhuis manifolds $(M_{1},\pi_{1},r_{1})$ and $(M_{2},\pi_{2},r_{2})$ induces, for every $n\in\mathbb{N}$, an isomorphism $H^{1}_{\pi_{1}^{n}}(M_{1})\cong H^{1}_{\pi_{2}^{n}}(M_{2})$ which maps the modular class $\left[X_{\pi_{1}^{n}}\right]$ of $M_{1}$ into the modular class $\left[X_{\pi_{2}^{n}}\right]$ of $M_{2}$.
Damianou and Fernandes [14] introduce a different notion of modular vector field, which is intrinsically associated with a Poisson-Nijenhuis manifold $(M,\pi,r)$. This modular vector field, denoted by $X_{r}$, is a Poisson vector field relative to the second Poisson structure of the hierarchy. When its associated Poisson cohomology class vanishes, $X_{r}$ becomes a bi-hamiltonian vector field. The modular vector field of the Poisson-Nijenhuis structure is defined using the modular vector fields of $\pi^{1}$ and $\pi$ relative to an arbitrary volume form $\nu$; specifically, $X_{r}\in\mathfrak{X}(M)$ is given by
$X_{r}=X_{\pi^{1}}^{\nu}-r\left(X_{\pi}^{\nu}\right).$ (57)
This vector field is independent of the choice of volume form $\nu$. Analogous to [12, Corollary 7], we expect the following to be true.
Proposition 4.23.

### v2

@@ 4.3. Global and infinitesimal Morita equivalences of compatible structures
is a Morita equivalence of the Dirac-Nijenhuis structures $\big(M_{i},L_{i}^{(0,n)},r_{i}\big)$, where $\big(\varpi^{(n)}\big)^{\flat}\coloneqq\varpi^{\flat}\circ J^{n}$.
We proceed by induction on $n$. For the base case $n=1$, we first show that the maps $\mu_{i}\colon\left(P,L_{\varpi^{(1)}}\right)\to\big(M_{i},L_{i}^{(0,1)}\big)$ are forward Dirac maps. Because $\mu_{i}\colon(P,L_{\varpi})\to(M_{i},L_{i})$ is already a forward Dirac map, elements in the fiber $(L_{i})_{\mu_{i}(p)}$ take the form $\zeta=(\mathrm{d}\mu_{i}(v),\alpha)$, where $v\in T_{p}P$ satisfies $\varpi^{\flat}(v)=(\mathrm{d}\mu_{i})^{*}\alpha$. By definition, elements in $L_{i}^{(0,1)}$ are of the form $(\mathrm{d}\mu_{i}(v),r_{i}^{*}\alpha)$. Because $(\varpi,J)$ is a compatible pair, we have $\big(\varpi^{(1)}\big)^{\flat}=\varpi^{\flat}\circ J=J^{*}\varpi^{\flat}$. Furthermore, the assumption that $J$ is $\mu_{i}$-related to $r_{i}$ implies $J^{*}\circ\mathrm{d}\mu_{i}^{*}=\mathrm{d}\mu_{i}^{*}\circ r_{i}^{*}$. Consequently, we obtain
$\big(\varpi^{(1)}\big)^{\flat}(v)=J^{*}\varpi^{\flat}(v)=J^{*}(\mathrm{d}\mu_{i})^{*}\alpha=(\mathrm{d}\mu_{i})^{*}(r_{i}^{*}\alpha).$
This establishes that $\mu_{i}\colon(P,L_{\varpi^{(1)}})\to\big(M_{i},L_{i}^{(0,1)}\big)$ is a forward Dirac map.
Second, we verify that the actions induced by these forward Dirac maps define an infinitesimal Morita bibundle. Denote by $\mathscr{a}_{i}\colon\Gamma(L_{i})\to\mathfrak{X}(P)$ the actions on $P$ induced by the Dirac structures $L_{i}$. For $i\in\{1,2\}$, the corresponding actions induced by the structures $L_{i}^{(0,1)}$, denoted by $\mathscr{a}_{i}^{(0,1)}\colon\Gamma\big(L_{i}^{(0,1)}\big)\to\mathfrak{X}(P)$, map sections of the form $\xi=(\mathrm{id}_{TM_{i}},r_{i}^{*})(\zeta)\in\Gamma\big(L_{i}^{(0,1)}\big)$—where $\zeta=(\mathrm{d}\mu_{i}(v),\alpha)\in\Gamma(L_{i})$—to the vector field $v$. In other words, we have
(55) $\mathscr{a}_{i}^{(0,1)}(\xi)=\mathscr{a}_{i}(\zeta).$
@@
Since the images of these actions coincide exactly with the vector fields of the original infinitesimal bibundle, it follows immediately that they satisfy all conditions for the Morita equivalence of both Dirac structures (Definition 4.11) and infinitesimal Nijenhuis structures (Definition 3.13). Furthermore, because the 22-form $\varpi^{(1)}$ is compatible with $J$, we have successfully constructed a Dirac–Nijenhuis Morita equivalence for the base case $n=1$.
Now, assume inductively that the result holds for a given integer $k\geq 1$; namely, that the structures $\big(M_{1},L_{1}^{(0,k)},r_{1}\big)$ and $\big(M_{2},L_{2}^{(0,k)},r_{2}\big)$ are Morita equivalent via the infinitesimal bibundle $\big(P,\varpi^{(k)},J\big)$. Applying a computation identical to that of the base case, we obtain
$\big(\varpi^{(k+1)}\big)^{\flat}(v)=(\mathrm{d}\mu_{i})^{*}\big((r_{i}^{*})^{k+1}\alpha\big)$
for $i\in\{1,2\}$. This identity implies that the maps $\mu_{i}\colon\left(P,L_{\varpi^{(k+1)}}\right)\to\big(M_{i},L_{i}^{(0,k+1)}\big)$ are forward Dirac maps, and that the actions of the Dirac structures $L_{i}^{(0,k+1)}$ satisfy
(56) $\mathscr{a}_{i}^{(0,k+1)}\big((\mathrm{id}_{TM_{i}},(r_{i}^{*})^{k+1})(\zeta)\big)=\mathscr{a}_{i}(\zeta).$
Because the images of these actions coincide with those of the original structures, and since $\varpi^{(k+1)}$ remains compatible with $J$ by construction, it follows that $M_{1}\xleftarrow{\ \mu_{1}\ }P\xrightarrow{\ \mu_{2}\ }M_{2}$ defines a Morita equivalence for the case $k+1$. This completes the induction and finishes the proof. ∎
@@
Let $(M_{1},\pi_{1},r_{1})$ and $(M_{2},\pi_{2},r_{2})$ be Morita equivalent Poisson–Nijenhuis structures via a bibundle
@@
If $J\in\Omega^{1}(P,TP)$ is an invertible Nijenhuis $(1,1)$-tensor field, then the Poisson–Nijenhuis structures $(M_{1},\pi_{1}^{(n)},r_{1})$ and $(M_{2},\pi_{2}^{(n)},r_{2})$ are (weakly) Morita equivalent for all $n\in\mathbb{N}$.
@@
To obtain a (strong) Morita equivalence, this proposition requires the assumption that the infinitesimal actions associated with the hierarchy are complete. These actions are defined by deforming the original actions $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ with powers of $J$ (as detailed in the construction below) and are not automatically complete. Without this completeness assumption, the structures are only weakly Morita equivalent in the sense of Ginzburg [undefs].
@@
Since $(\varpi,J)$ is a compatible pair and $\varpi$ is symplectic, $(P,\pi_{\varpi},J)$ is a Poisson-Nijenhuis manifold, where $\pi_{\varpi}=-(\varpi^{\flat})^{-1}$. Consequently, there exists a hierarchy of Poisson-Nijenhuis structures $(\pi_{\varpi}^{(n)},J)$ on $P$. Note that each $\pi_{\varpi}^{(n)}$ is a non-degenerate Poisson structure. Indeed, $J^{n}$ is invertible since $J$ is invertible; furthermore, $\pi_{\varpi}^{\sharp}$ is an isomorphism by the non-degeneracy of $\varpi$. It follows that the composition $\big(\pi_{\varpi}^{(n)}\big)^{\sharp}=J^{n}\circ\pi_{\varpi}^{\sharp}$ is invertible. We define the symplectic forms $\varpi^{(n)}\in\Omega^{2}(P)$ by $\big(\varpi^{(n)}\big)^{\flat}\coloneqq-\big(\pi_{\varpi}^{(n)\sharp}\big)^{-1}$, i.e. $\big(\varpi^{(n)}\big)^{\flat}=\varpi^{\flat}\circ J^{-n}$. Finally, the bibundle for the $n$-th level of the hierarchy is given by the bibundle
$\left(M_{1},\pi_{1}^{(n)},r_{1}\right)\xleftarrow{\quad\mu_{1}\quad}\left(P,\varpi^{(n)},J\right)\xrightarrow{\quad\mu_{2}\quad}\left(M_{2},\pi_{2}^{(n)},r_{2}\right).$
The infinitesimal actions are defined by deforming the original actions via the tensor field $J^{n}$, which yields the natural actions on the corresponding symplectic realizations of the Poisson manifolds. Let $\mathscr{a}^{(n)}_{i}\colon\Gamma(T^{*}M_{i})\to\mathfrak{X}(P)$ denote the actions of the cotangent algebroids associated to the Poisson manifolds $\big(M_{i},\pi_{i}^{(n)}\big)$. For any $\alpha\in\Omega^{1}(M_{1})$ and $\beta\in\Omega^{1}(M_{2})$, these actions are explicitly given by
(57) $\mathscr{a}^{(n)}_{1}(\alpha)=\big(\pi_{\varpi}^{(n)}\big)^{\sharp}(\mu_{1}^{*}\alpha)\qquad\text{and}\qquad\mathscr{a}^{(n)}_{2}(\beta)=\big(\pi^{(n)}_{\varpi}\big)^{\sharp}(\mu_{2}^{*}\beta).$
Provided the completeness of these infinitesimal actions, establishing Morita equivalence at the $n$-th level reduces to verifying the remaining conditions for an infinitesimal Morita bibundle. We now proceed to verify these conditions.
The commutativity of the actions $\mathscr{a}_{1}^{(n)}$ and $\mathscr{a}_{2}^{(n)}$ follows directly from the compatibility of the pair $(\pi_{\varpi},J)$ and the assumption that $J$ is $\mu_{i}$-related to $r_{i}$ for each $i\in\{1,2\}$. Specifically, we compute
$\left[\mathscr{a}^{(n)}_{1}(\alpha),\mathscr{a}^{(n)}_{2}(\beta)\right]$ $=\left[J^{n}\big(\pi_{\varpi}^{\sharp}(\mu_{1}^{*}\alpha)\big),J^{n}\big(\pi_{\varpi}^{\sharp}(\mu_{2}^{*}\beta)\big)\right]$
$=\left[\pi_{\varpi}^{\sharp}\left(\mu_{1}^{*}\left((r_{1}^{*})^{n}\alpha\right)\right),\pi_{\varpi}^{\sharp}\left(\mu_{2}^{*}\left((r_{2}^{*})^{n}\beta\right)\right)\right]$
$=0.$
Finally, the non-degeneracy of $J$ guarantees that for every $p\in P$, the linear maps $\mathscr{a}_{1\,p}^{(n)}\colon(\mu_{1}^{*}(T^{*}M_{1}))_{p}\to T_{p}P$ and $\mathscr{a}_{2\,p}^{(n)}\colon(\mu_{2}^{*}(T^{*}M_{2}))_{p}\to T_{p}P$ map injectively onto $\operatorname{ker}\mathrm{d}\mu_{2\,p}$ and $\operatorname{ker}\mathrm{d}\mu_{1\,p}$, respectively.
Remark 4.23 (Non-degenerate infinitesimal Morita bibundle).
The framework of Proposition 4.21 is most naturally framed in terms of an infinitesimal bibundle endowed with a non-degenerate Poisson bivector and an invertible, compatible Nijenhuis tensor field. Specifically, the infinitesimal Morita equivalence of Poisson–Nijenhuis manifolds is given by
$(M_{1},\pi_{1},r_{1})\xleftarrow{\quad\mu_{1}\quad}(P,\pi_{\varpi},J)\xrightarrow{\quad\mu_{2}\quad}(M_{2},\pi_{2},r_{2}).$
In this setting, the triple $(r_{1},J,r_{2})$ generates the hierarchy of Morita equivalences of Poisson manifolds, yielding the following diagram:
$\left(M_{1},\pi_{1}\right)$$\left(M_{1},\pi_{1}^{(1)}\right)$$\cdots$$\left(M_{1},\pi_{1}^{(n)}\right)$$\cdots$$\left(P,\pi_{\varpi}\right)$$\left(P,\pi_{\varpi}^{(1)}\right)$$\cdots$$\left(P,\pi_{\varpi}^{(n)}\right)$$\cdots$$\left(M_{2},\pi_{2}\right)$$\left(M_{2},\pi_{2}^{(1)}\right)$$\cdots$$\left(M_{2},\pi_{2}^{(n)}\right)$$\cdots$$\scriptstyle{r_{1}}$$\scriptstyle{r_{1}}$$\scriptstyle{r_{1}}$$\scriptstyle{r_{1}}$$\scriptstyle{\mu_{1}}$$\scriptstyle{\mu_{2}}$$\scriptstyle{J}$$\scriptstyle{\mu_{1}}$$\scriptstyle{\mu_{2}}$$\scriptstyle{J}$$\scriptstyle{J}$$\scriptstyle{\mu_{1}}$$\scriptstyle{\mu_{2}}$$\scriptstyle{J}$$\scriptstyle{r_{2}}$$\scriptstyle{r_{2}}$$\scriptstyle{r_{2}}$$\scriptstyle{r_{2}}$
We call an infinitesimal bibundle with this structure a non-degenerate Morita equivalence of Poisson–Nijenhuis manifolds.
4.4.1. Modular vector field of a Poisson-Nijenhuis manifold.
@@
We now comment on the cohomological consequences arising from the hierarchy of Morita equivalent Poisson–Nijenhuis structures. Henceforth, we consider a non-degenerate infinitesimal Morita bibundle between Poisson–Nijenhuis structures $(M_{1},\pi_{1},r_{1})$ and $(M_{2},\pi_{2},r_{2})$. By Proposition 4.21, there exists a corresponding hierarchy of infinitesimal bibundles
$\left(M_{1},\pi_{1}^{(n)},r_{1}\right)\xleftarrow{\quad\mu_{1}\quad}\left(P,\pi_{\varpi}^{(n)},J\right)\xrightarrow{\quad\mu_{2}\quad}\left(M_{2},\pi_{2}^{(n)},r_{2}\right),$
which constitute weak Morita equivalences (as discussed in Remark 4.22). The results of the present subsection remain valid in this weaker setting.
@@
Every Poisson manifold $(M,\pi)$ induces a differential $\mathrm{d}_{\pi}\colon\mathfrak{X}^{k}(M)\to\mathfrak{X}^{k+1}(M)$. The cohomology $H^{\bullet}_{\pi}(M)$ of the complex $(\mathfrak{X}^{\bullet}(M),\mathrm{d}_{\pi})$ is called the Poisson cohomology of $(M,\pi)$. It was shown in [undeft] that Morita equivalent Poisson manifolds $(M_{1},\pi_{1})$ and $(M_{2},\pi_{2})$ have isomorphic first Poisson cohomology groups; that is, $H^{1}_{\pi_{1}}(M_{1})\cong H^{1}_{\pi_{2}}(M_{2})$. In our setting, we thus obtain a hierarchy of isomorphic first Poisson cohomology groups. To simplify notation, we write $H^{1}_{(n)}(M_{i})\coloneqq H^{1}_{\pi_{i}^{(n)}}(M_{i})$. For every $n\in\mathbb{N}$, it then follows that
$H^{1}_{(n)}(M_{1})\cong H^{1}_{(n)}(M_{2}).$
@@
We recall the modular vector field of a Poisson manifold $(M,\pi)$. Suppose initially that $M$ is orientable. The divergence of a vector field $v\in\mathfrak{X}(M)$ with respect to a volume form $\nu$ is the unique function $\mathrm{div}_{\nu}(v)\in C^{\infty}(M)$ such that $\mathscr{L}_{v}(\nu)=\mathrm{div}_{\nu}(v)\nu$. For any function $f\in C^{\infty}(M)$, let $X_{f}=\pi^{\sharp}(\mathrm{d}f)$ denote its Hamiltonian vector field. The modular vector field of $(M,\pi)$ with respect to $\nu$ is the Poisson vector field $\mathscr{X}^{\nu}\in\mathfrak{X}(M)$ defined by
$\mathscr{X}_{\nu}(f)\coloneqq\mathrm{div}_{\nu}\left(X_{f}\right).$
@@
While this vector field depends on the choice of the volume form, its cohomology class $\left[\mathscr{X}_{\nu}\right]\in H^{1}_{\pi}(M)$ does not. We therefore write $\mathscr{X}]\coloneqq[\mathscr{X}_{\nu}]$, and this class is called the modular class of the Poisson manifold $(M,\pi)$. By replacing volume forms with smooth densities in the definition of the modular vector field, we may drop the orientability requirement on the manifold $M$.